# Topic-Based Watermarks for Large Language Models

### 02 Mar 2026 v5

@@ C.1 Strict and Sliding Window Detection
As discussed in §5.4, both the strict and sliding window detection schemes exhibit degraded performance due to imperfect topic matching between input prompts and output text. This limitation arises because the detection pipeline assumes alignment between topics extracted from the input and the generated output, an assumption that often fails in general-purpose, black-box deployment. However, there are domains where such access to the input text is natural and the tradeoff becomes favorable.
@@ E.4 Efficiency Results
Figures 5 and 6 present generation times for OPT-6.7B and OPT-2.7B, respectively. We measure the efficiency overhead introduced by each watermarking method, using 10 samples from C4 and generating sequences of lengths {100, 200, 300, 400, 500}. For each token-length setting, we record the average generation time over the 10 samples.
TBW introduces negligible overhead relative to non-watermarked generation across all sequence lengths, matching lightweight methods (KGW, SynthID, Unigram). In contrast, EXP-Edit requires multiple re-ranking passes and SIR incurs additional complexity, resulting in noticeably higher generation times. These trends hold consistently across both model scales, confirming TBW’s practicality for deployment.
@@ E.5 Comparison to ITS-Edit
E.6 TBW Robustness ROC Curves
@@
E.7 False Positive Rate (FPR) on Human-Written Text
E.7.1 Watermarking Methods FPR
@@ E.8 Detection Attacks
For completeness, we evaluate our method against non-watermarked text using two classes of detection strategies: (i) statistical hypothesis tests over generated output and (ii) supervised machine learning classification trained to separate watermarked and non-watermarked corpora. We conduct experiments across our two representative models (Gemma-7B and OPT-6.7B) and evaluate at bias strengths $\delta=2.0$ and $\delta=3.0$. These values match those used in our main evaluations of text quality and robustness, allowing us to study detection performance under the same operating points and to compare tradeoffs across quality, robustness, and detectability. While these experiments provide additional perspective, they rely on assumptions such as access to large paired corpora of watermarked and non-watermarked outputs that rarely hold in deployed systems, and thus fall outside the scope of practical adversaries targeted by our primary threat model.
E.8.1 Statistical Tests
@@
We note that for Gemma, the watermarked outputs exhibit slightly flatter token distributions compared to the corresponding non-watermarked outputs. While this effect is subtle and not readily visible without corpus-level aggregation, it is consistent across bias strengths and may indicate a mild redistribution of lexical mass. Nonetheless, in realistic deployments, attackers would not have access to paired corpora of watermarked and non-watermarked text from the same model, and thus could not easily identify subtle differences even if they existed. The closest analogue is Google’s evaluation of SynthID-Text, where generations from the same model included both watermarked and non-watermarked outputs without disclosure to users SythID-Team (2024); Dathathri et al. (2024). However, since TBW preserves the same dominant distribution of function words across conditions, such comparisons likewise fail to yield clear evidence of watermark presence. In short, statistical tests do not provide a practical detection vector against TBW.
E.8.2 Machine Learning Classification
@@ E.8.3 Discussion and Implications
E.9 Maximum $z$-Score Detection: Detailed Analysis
@@ E.9.1 Performance Metrics
E.9.2 $z$-Score Distribution Analysis
@@ F.1 TBW Strength Sensitivity
Detection-Quality Trade-off. Figures 12 and 12 illustrate the core trade-off in watermark strength. As $\delta$ increases, detection scores rise monotonically across all $\tau$ values, reaching $z$-scores above 24 at $\delta=10.0$ (Figure 12). However, higher strengths gradually reduce lexical diversity (Figure 12), with the most notable changes occurring between $\delta=2.0$ and $\delta=5.0$. This pattern underscores the importance of moderate watermark strengths for balancing detectability and output quality. Notably, $\tau$ has minimal influence on either detection performance or diversity, indicating that semantic alignment can be optimized independently of other quality metrics. Additional visualizations covering the full parameter space are provided in F.2.
@@
F.2 Parameter Space Analysis
Figures 13 and 14 provide complementary views of the complete parameter space. The scatter plots in Figure 13 reveal that as detection scores increase with higher $\delta$ values, lexical diversity consistently decreases across all Distinct-N metrics, with threshold values showing negligible variation. This relationship is further confirmed by the heat map in Figure 14, which illustrates that lower bias strengths ($\delta\leq 2.0$) maintain high Distinct-N scores (>0.7>0.7) across all thresholds, while aggressive watermarking ($\delta\geq 5.0$) results in substantial quality degradation (Distinct-N <0.5<0.5). The vertical consistency in the heat map reinforces that similarity threshold $\tau$ can be selected based on semantic considerations without significantly affecting lexical diversity.
@@ F.4 Scaling the Number of Topics
In our main experiments we instantiated topic-based watermarking (TBW) with $K=4$ topic lists. While this setting is sufficient in many scenarios, model owners may wish to scale $K$ to better cover niche domains or to improve generalization. Under our partitioning scheme, increasing $K$ reduces each topic list’s share of the vocabulary, which can weaken the per-token watermark signal (since the logit bias applies to a smaller token subset) and thereby affect detection and text quality. We therefore ablate $K\in\{4,8,16,32\}$.
@@ F.4.2 Results
Detection strength vs. $K$. We evaluate the detector’s $z$-score (higher is stronger and more resilient to edits and paraphrases). Figure 15 reports the mean $z$-score under maximum-$z$ as a function of $K$. As expected, the signal decreases as $K$ grows, yet remains strong: from $\sim 11$ at $K=4$ down to $\sim 7$ at $K=32$. Notably, even at $K=32$ the $z$-scores are comparable to, if slightly below, Unigram and KGW baselines at the same $\delta$, despite the much smaller per-list partitions.
[…]
@@
[…]
Number of Topics ($K$) Detection Time (s)
4 $2.002\pm 0.041$
8 $2.910\pm 0.047$
16 $3.833\pm 0.046$
32 $6.565\pm 0.066$
@@ Appendix H Watermarked Text Examples
idx Model Prompt Real Completion No Watermark Watermarked

### 15 Apr 2026 v6

@@ C.1 Strict and Sliding Window Detection
As discussed in §5.4, both the strict and sliding window detection schemes exhibit degraded performance due to imperfect topic matching between input prompts and output text. This limitation arises because the detection pipeline assumes alignment between topics extracted from the input and the generated output, an assumption that often fails in general-purpose, black-box deployment. However, there are domains where such access to the input text is natural and the trade-off becomes favorable.
@@ E.4 Efficiency Results
To verify that TBW’s efficiency advantage holds across model scales, we repeat the generation time evaluation from §5.5 on OPT-2.7B. Figure 7 confirms that the same trends observed on OPT-6.7B hold at smaller scale since TBW introduces negligible overhead relative to non-watermarked generation, matching lightweight methods across all sequence lengths.
@@ E.5 Comparison to ITS-Edit
Method Time PPL No Attack DIPPER
(s/sample) (mean $\pm$ std) ROC-AUC TPR@1% ROC-AUC TPR@1%
TBW ($\delta$=2.0) 5.23 $\pm$ 1.93 12.36 $\pm$ 3.64 0.971 0.879 0.694 0.222
TBW ($\delta$=3.0) 5.08 $\pm$ 2.01 12.46 $\pm$ 4.80 0.986 0.949 0.802 0.357
k-SemStamp 65.93 $\pm$ 75.03 9.16 $\pm$ 2.85 0.960 0.750 0.743 0.150
PostMark 29.61 $\pm$ 17.56 12.43 $\pm$ 2.90 0.968 0.920 0.780 0.210
Table 10: Comparison of TBW against semantic black-box watermarking methods on OPT-6.7B over 100 C4 prompts. Bold indicates best results. TPR@1% denotes TPR at 1% FPR.
@@
E.6 Comparison to Semantic Watermarks
To situate TBW relative to semantic watermarking approaches that operate via post-processing, we compare against PostMark Chang et al. (2024) and k-SemStamp Hou et al. (2024b). A key methodological distinction is that both PostMark and k-SemStamp embed watermarks by rewriting already-generated text through additional LLM inference passes, whereas TBW embeds the watermark during a single standard decoding pass. This difference has direct implications for generation cost and deployment complexity.
We evaluate all methods over 100 C4 prompts on OPT-6.7B, measuring perplexity, generation time, ROC-AUC, and TPR@1%FPR under both no-attack and DIPPER paraphrasing conditions. TBW is evaluated at two watermark strengths ($\delta=2.0$ and $\delta=3.0$) to illustrate the quality-robustness trade-off. Results are reported in Table 10.
[…]
E.7 TBW Robustness ROC Curves
@@
E.8 False Positive Rate (FPR) on Human-Written Text
E.8.1 Watermarking Methods FPR
@@ E.9 Detection Attacks
For completeness, we evaluate our method against non-watermarked text using two classes of detection strategies: (i) statistical hypothesis tests over generated output and (ii) supervised machine learning classification trained to separate watermarked and non-watermarked corpora. We conduct experiments across our two representative models (Gemma-7B and OPT-6.7B) and evaluate at bias strengths $\delta=2.0$ and $\delta=3.0$. These values match those used in our main evaluations of text quality and robustness, allowing us to study detection performance under the same operating points and to compare trade-offs across quality, robustness, and detectability. While these experiments provide additional perspective, they rely on assumptions such as access to large paired corpora of watermarked and non-watermarked outputs that rarely hold in deployed systems, and thus fall outside the scope of practical adversaries targeted by our primary threat model.
E.9.1 Statistical Tests
@@
We note that for Gemma, the watermarked outputs exhibit slightly flatter token distributions compared to the corresponding non-watermarked outputs. While this effect is subtle and not readily visible without corpus-level aggregation, it is consistent across bias strengths and may indicate a mild redistribution of lexical mass. Nonetheless, in realistic deployments, attackers would not have access to paired corpora of watermarked and non-watermarked text from the same model, and thus could not easily identify subtle differences even if they existed. The closest analogue is Google’s evaluation of SynthID-Text, where generations from the same model included both watermarked and non-watermarked outputs without disclosure to users SynthID-Team (2024); Dathathri et al. (2024). However, since TBW preserves the same dominant distribution of function words across conditions, such comparisons likewise fail to yield clear evidence of watermark presence. In short, statistical tests do not provide a practical detection vector against TBW.
E.9.2 Machine Learning Classification
@@ E.9.3 Discussion and Implications
E.10 Maximum $z$-Score Detection: Detailed Analysis
@@ E.10.1 Performance Metrics
E.10.2 $z$-Score Distribution Analysis
@@ F.1 TBW Strength Sensitivity
Detection-Quality Trade-off. Figures 13 and 13 illustrate the core trade-off in watermark strength. As $\delta$ increases, detection scores rise monotonically across all $\tau$ values, reaching $z$-scores above 24 at $\delta=10.0$ (Figure 13). However, higher strengths gradually reduce lexical diversity (Figure 13), with the most notable changes occurring between $\delta=2.0$ and $\delta=5.0$. This pattern underscores the importance of moderate watermark strengths for balancing detectability and output quality. Notably, $\tau$ has minimal influence on either detection performance or diversity, indicating that semantic alignment can be optimized independently of other quality metrics. Additional visualizations covering the full parameter space are provided in F.2.
@@
F.2 Parameter Space Analysis
Figures 14 and 15 provide complementary views of the complete parameter space. The scatter plots in Figure 14 reveal that as detection scores increase with higher $\delta$ values, lexical diversity consistently decreases across all Distinct-N metrics, with threshold values showing negligible variation. This relationship is further confirmed by the heat map in Figure 15, which illustrates that lower bias strengths ($\delta\leq 2.0$) maintain high Distinct-N scores (>0.7>0.7) across all thresholds, while aggressive watermarking ($\delta\geq 5.0$) results in substantial quality degradation (Distinct-N <0.5<0.5). The vertical consistency in the heat map reinforces that similarity threshold $\tau$ can be selected based on semantic considerations without significantly affecting lexical diversity.
@@ F.4 Scaling the Number of Topics
This section provides full experimental details and additional visualizations for the topic scalability analysis presented in §5.6.
@@ F.4.2 Results
idx Model Prompt Real Completion No Watermark Watermarked